# Universal Kerr effect in A-type antiferromagnets

### v1

@@ Abstract
Magneto-optic Kerr effect (MOKE) is a powerful probe of broken time-reversal symmetry ($\mathcal{T}$), typically used to study ferromagnets. While MOKE has been observed in some antiferromagnets (AFMs) with vanishing magnetization, it is often associated with structures whose symmetry is lower than basic collinear, bipartite order. In contrast, theory predicts a mechanism for MOKE intrinsic to all AFMs of A-type, i.e. layered AFMs in which ferromagnetic layers are antiferromagnetically aligned. Here we report the first experimental confirmation of this mechanism. We achieve this by measuring the imaginary component of MOKE as a function of photon energy in MnBi2Te4, an A-type AFM where $\mathcal{T}$ is preserved in combination with a translation. By comparing the experimental results with model calculations, we demonstrate that observable MOKE should be expected in all collinear A-type AFMs with out-of-plane spin order, thus enabling optical detection of AFM domains and expanding the scope of MOKE to few-layer AFMs.
@@
Figure 1: Schematics of representative magnetic structures. (a) Ferromagnet, where $\mathcal{T}$ reverses the sign of the order parameter; (b) A-type antiferromagnet, where the sign reversal by $\mathcal{T}$ can be undone by a translation by half a magnetic unit cell ($S_{1/2}$); (c) Thin flakes of an A-type antiferromagnet: flakes of odd layer number $N$ have a net magnetic moment, while those with even $N$ do not. (d) Two layers of the MnBi2Te4  crystal and magnetic structure, demonstrating the A-type AFM phase. Opposite spins are related by $S_{1/2}$.
@@
We investigate MnBi2Te4, a layered topological AFM with a Néel temperature of 25 K [19], as an example of a $\mathcal{T}S_{1/2}$ symmetric AFM. Spins in individual Mn layers in MnBi2Te4  are parallel to each other, and antiparallel to the spins in neighboring layers (Fig. 1d). This order, called A-type, persists to the surface of MnBi2Te4 [20]. MOKE was recently reported in the wavelength range of $500-1000\,\mathrm{nm}$ in thin flakes of MnBi2Te4, in samples with an odd number of layers ($N$), which have a net magnetization, and in samples with an even $N$, which do not [21, 22, 23, 24]. While symmetry permits MOKE for all $N$, it does not identify the underlying mechanism.
@@
A distinct mechanism for MOKE in $\mathcal{T}S_{1/2}$ symmetric AFMs was proposed by Dzyaloshinskii and Papamichail [31]. They treated the layered AFM as a stack of FMs with alternating spin direction. The dielectric tensor of each layer is assumed to be the same as in an unbounded medium with the same FM order, and MOKE arises due to the variation in dielectric properties as the wave propagates. In this ‘alternating FM’ approach the surface conductance does not play a special role, in contrast to the axion electrodynamics mechanism. To the best of our knowledge, MOKE arising from the alternating FM effect has not been experimentally recognized prior to this work.
Below we report measurements of MOKE in ‘infinite’ ($\sim 136\,\mathrm{nm}$ thick, Fig. S1) layer MnBi2Te4 as a function of magnetic field, $H$, and wavelength of the optical probe, $\lambda$. We find that non-zero MOKE appears in the $M=0$ state, and that its spectrum is quantitatively captured by the alternating FM model [31], without invoking axion electrodynamics. The alternating FM mechanism is allowed in all collinear AFMs whose spins are oriented perpendicular to the layers, regardless of the number of layers, band topology or symmetry, opening the door for MOKE studies of a broader class of AFMs than was previously considered feasible.
@@
Figure 2: Magneto-optical measurements of MnBi2Te4. (a) Reflection circular dichroism (RCD) measured at $525\,\mathrm{nm}$ as a function of magnetic field, showing a jump at the spin-flop transition ($H_{SF}\approx 3.4\,\mathrm{T}$) and a hysteresis loop with a coercive field of $H_{C}\approx 1\,\mathrm{T}$. A small setup-induced background contribution has been subtracted (Fig. S5). RCD is compared with bulk magnetization measurements taken at $15\,\mathrm{K}$ with the field applied along the crystallographic $c$-axis (black line). The inset illustrates the bulk spin-flop transition. (b) Spatially resolved RCD measurements at $H=0$, revealing AFM domain structures that change between cooldowns.
The complex MOKE angle is defined by the relation, $\Theta\equiv-i\delta r/r$, where the reflectivities of incident LCP and RCP light are given by $r=r_{0}\pm\delta r$ (the sign and polarization conventions are defined in the Methods section). Throughout this paper we show measurements of the imaginary part of $\Theta$, which is known as reflection circular dichroism (RCD). We measure RCD using a photoelastic modulator to vary the helicity of the incident light at $50\,\mathrm{kHz}$, using a setup illustrated in Fig. S2. The change in reflectivity synchronous with the helicity modulation is proportional to RCD (see Methods for details of the experiment).
@@
The existence of two discontinuities in RCD vs. $H$ provides an unusual opportunity to compare the MOKE spectrum in a phase with $M\neq 0$, that is the spin-flop (SF) phase, and a $\mathcal{T}S_{1/2}$symmetric AFM ($M=0$) phase in the same material. We measured the magnetic field dependence of RCD in the photon energy range from $1.34\,\mathrm{eV}$ ($925\,\mathrm{nm}$) to $3.22\,\mathrm{eV}$ ($385\,\mathrm{nm}$) (Fig. S6), and extracted RCD spectra in the two phases. In Fig. 3a we compare the RCD spectrum corresponding to the SF transition with the RCD spectrum in the AFM phase. The AFM spectrum is half of the difference between RCD at $H=0^{+}$ and $H=0^{-}$, where $H=0^{+}$ and $H=0^{-}$ denote zero field as approached from positive and negative fields. The SF spectrum, corresponding to the jump in magnetization across the transition, is given by $\Theta_{SF}=\Theta(3.4\,\mathrm{T})-\left[\Theta(3.2\,\mathrm{T})-\Theta(0^{+})\right]$, because $\Theta(3.2\,\mathrm{T})$ is a sum of the antiferromagnetic and magnetization contributions, while in the spin flop phase, at $3.4\,\mathrm{T}$, there is no antiferromagnetic contribution.
@@
Figure 3: Spectroscopic signature of MOKE in antiferromagnetic MnBi2Te4. (a) Measured RCD spectrum at the spin-flop transition (purple) and at zero field (red), showing distinct spectral shapes. For each photon energy, we measured the magnetic field dependence of RCD (Fig. S6). The AFM spectrum is half of the difference between RCD on the downward ($H=0^{+}$) and upward ($H=0^{-}$) field sweep, averaged across the hysteresis loop, between ±$\pm 0.9\,\mathrm{T}$. The SF spectrum is given by $\Theta_{SF}=\Theta(3.4\,\mathrm{T})-\left[\Theta(3.2\,\mathrm{T})-\Theta(0^{+})\right]$. The error bars for the $H=0^{+}$ measurement represent the standard deviation across this field range. The larger error bars for the SF spectrum reflect the fact fewer measurements were used to extract the corresponding RCD value (at $H=3.4\,\mathrm{T}$ and $H=3.2\,\mathrm{T}$). (b) RCD spectra computed from the Lorentz model (Eq. 1, with parameters from Table S1) for a ferromagnet (purple, schematic in panel c) and two antiferromagnetic domains (red and blue, schematics in panel d). The ferromagnetic RCD is scaled by a factor of $0.4$ from that calculated for a saturated moment, to be directly comparable to the measured RCD jump at the spin-flop transition.
@@
where $f$, $\omega_{0}$, and $\gamma$ are oscillator strength, resonant frequency, and damping. These parameters, together with a background contribution, $\varepsilon_{\infty}$, are chosen to reproduce the measured optical conductivity in the spectral range of interest, reported in Ref. [33]. The good agreement between the measured quantities and those calculated from the Lorentz model is shown in Fig. S3. Finally, $\delta\omega$, the difference in resonance frequency for LC and RC polarized light, is chosen to fit the magnitude of the measured RCD. The purple curve in Fig. 3b shows that the single oscillator model captures the main features of the observed ferromagnetic RCD spectrum (Fig. 3a).
Our quantitative description of RCD in the ferromagnetic phase of MnBi2Te4 leads us to consider the alternating FM model for the RCD in its AFM state: a stack of FM layers in which the sign of $\delta\omega$ alternates in neighboring layers. We use the transfer matrix formalism [34], following the approach in Ref. [35], to calculate the reflection coefficient from the stack. The number of layers is 100, and the thickness of each layer, $d=\mathrm{1.3}{nm}$, corresponds to the separation between spins in MnBi2Te4. The simulated RCD spectra for the two $\mathcal{T}$-reversed AFM domains (Fig. 3d), are shown in red and blue in Fig. 3b. Remarkably, this minimal model captures the spectral shape and the magnitude of RCD in the AFM phase of MnBi2Te4, without introducing free parameters beyond those that describe the FM phase. In the SI we extend this analysis by using a dielectric function computed from a density functional theory (DFT) calculation to construct the alternating FM model, and we find that this approach also yields the correct magnitude of RCD in the AFM phase (Fig. S4). To the best of our knowledge, this is the first experimental evidence for an RCD arising from an alternating FM model in an AFM.
@@
Further, the dichroism spectra for even and odd layer stacks are strikingly different (Fig. 4c, d). For $N=8$, the RCD spectrum is identical to the spectrum of the bulk AFM crystal (Fig. 3b), while spectrum for $N=9$ resembles that of the bulk SF phase, as $M\neq 0$ for odd layer stacks. The maximum RCD is about three times larger for $N=9$ than for $N=8$. However, these calculations assume that samples are suspended in vacuum, which is not a realistic experimental scenario. The transfer matrix approach allows us to include the substrate in the calculation of reflectivity [35], and we find that doing so preferentially suppresses dichroism in odd-layer samples. In Fig. S7 we show that the maximum magnitude of RCD for samples on sapphire and diamond substrates, used in Ref. [23], is approximately equal in magnitude for even and odd layer stacks, while they maintain the characteristic spectral shapes shown in Fig. 4c. We therefore find that the alternating FM model with realistic parameters captures the main experimental results on $N$-even and $N$-odd flakes of MnBi2Te4 reported in [23], as well as on ‘infinite’ layer flakes reported here.
@@
Eq. 2 showcases the common microscopic origin of the MOKE response of an AFM bilayer and that of the corresponding ferromagnet, but it also reveals a surprising aspect of the link between them. It is natural to assume that a $\mathcal{T}S_{1/2}$ symmetric type-A AFM can exhibit nonzero MOKE only if strong absorption (governed by the imaginary part of $n$) leads to disproportionate sampling of the top layer. However, Eq. 2 shows that this is not correct: the magnitude of $\Theta_{AFM}$ depends on the absolute value of the index, $|n|$, and is nonzero even in the absence of dissipation. The phase of $n$ appears in the phase shift between $\Theta_{AFM}$ and $\Theta_{FM}$, which manifests as the difference in the spectrum of RCD between AFM and FM structures, as seen for example in Fig. 3a.
Figure 5: Illustration of the model for MOKE in an antiferromagnetic bilayer. (a) Schematic representation of a bilayer, consisting of two layers of indices of refraction $n_{1,2}=n\pm\delta n$ when probed with circularly polarized light. $\delta n$ is induced by the opposite magnetization in the two layers, ±$\pm M_{z}$. (b,c) Simplified representation of a bilayer, where the layers are replaced by two infinitesimally thin sheets of conductances $G_{1,2}=G_{0}\pm\delta G$, separated by a dielectric medium of thickness $d$ and index of refraction $n$. The propagation-induced differences between fields emitted from the two sheets cancel in (b) transmission, but not in (c) reflection, leading to vanishing TCD and non-zero RCD.
@@
$E_{1,2}(z)=\frac{\mu_{0}c}{2}K_{1,2}e^{ik\left|z\pm d/2\right|},$ (4)
@@
$E_{T}(z)=\mu_{0}cE_{0}\left(\frac{G_{1}+G_{2}}{2}\right)e^{ikz}=\mu_{0}cE_{0}G_{0}e^{ikz}.$ (5)
@@
In contrast, the propagation factors for fields radiated in the $-z$ direction add, rather than cancel: the incident EM wave travels further to excite current in layer 2 and its radiated field propagates further before reaching an observer at $z<-d/2$ (Fig. 5c). The total reflected electric field is given by,
@@ Conclusion
In this paper we identify alternating FM layers as the origin of MOKE in an A-type antiferromagnet with $\mathcal{T}S_{1/2}$ symmetry, therefore providing first experimental confirmation of the mechanism proposed by Dzyaloshinskii and Papamichail [31]. Within this theory, $\Theta_{AFM}$ is entirely determined by the MOKE response of the ferromagnetic layers that are the building blocks of the AFM, allowing us to test it directly by tuning MnBi2Te4  across the spin-flop transition, thus accessing $\Theta_{AFM}$ and $\Theta_{FM}$ in the same crystal. Based on the quantitative agreement between the measured RCD spectrum in the AFM phase and the theoretical prediction based solely on the FM spectrum, we conclude that MOKE observed at visible and near infrared wavelengths in MnBi2Te4 arises from alternating FM layers, rather than axion electrodynamics associated with band topology. Axion electrodynamics may still dominate the MOKE response at terahertz frequencies below the bulk band gap [27, 25, 39, 40], where the topological effects are enhanced, and there are no other optical transitions that could overwhelm their contribution.
The effect we observe is not limited to MnBi2Te4, and is generically present in A-type antiferromagnets with spins out of plane. We believe that it has not been recognized experimentally in part because the signal vanishes when averaged over atomic steps, naturally present in as-grown bulk crystals. However, optical experiments in van der Waals materials regularly probe surfaces that are atomically flat across the length scale of a laser beam. We note that some previous sightings of MOKE in van der Waals AFMs with an even number of layers were conditionally attributed to extrinsic magnetization [41, 42, 22], given the widespread conviction that intrinsic MOKE was forbidden by symmetry. While this explanation is certainly plausible, our findings show that intrinsic MOKE in A-type AFMs is always allowed, and other mechanisms should be considered when $\Theta_{AFM}$ cannot account for the magnitude, or the spectrum, of the observed signal.
@@
Our findings show that MOKE is a powerful and general tool for investigating $\mathcal{T}S_{1/2}$ symmetric AFMs. In this work we used it to image time-reversed domains, as shown for in Fig. 2b. Further, the nonzero $\Theta_{AFM}$ enabled the discovery that a magnetic field can select antiferromagnetic domains (Fig. 2a) in such magnets, challenging the conventional view that antiferromagnets cannot be controlled by external magnetic fields. Uncovering the mechanism behind field control is beyond the scope of this study, but we note that the weakened exchange interaction at the surface is a plausible cause [20, 43]. In summary, our results address fundamental questions concerning the origin of MOKE in AFMs, provide a strategy for studying and controlling their domains, and elucidate results on exfoliated van der Waals AFMs.

### v2

# Magneto-optical Kerr effect in an A-type antiferromagnet

@@ Abstract
Magneto-optic Kerr effect (MOKE) is a powerful probe of broken time-reversal symmetry ($\mathcal{T}$), typically used to study ferromagnets. While MOKE has been observed in some antiferromagnets (AFMs) with vanishing magnetization, it is often associated with structures whose symmetry is lower than basic collinear, bipartite order. In contrast, theory predicts a mechanism for MOKE intrinsic to all AFMs of A-type, i.e. layered AFMs in which ferromagnetic layers are antiferromagnetically aligned. Here we report the first experimental confirmation of this mechanism in a bulk AFM. We achieve this by measuring the imaginary component of MOKE as a function of photon energy in MnBi2Te4, an A-type AFM where $\mathcal{T}$ is preserved in combination with a translation. By comparing the experimental results with model calculations, we demonstrate that observable MOKE should be expected in all collinear A-type AFMs with out-of-plane spin order, thus enabling optical detection of AFM domains and expanding the scope of MOKE to few-layer AFMs.
@@
Figure 1: Schematics of representative magnetic structures. (a) Ferromagnet, where $\mathcal{T}$ reverses the sign of the order parameter; (b) A-type antiferromagnet, where the sign reversal by $\mathcal{T}$ can be undone by a translation by half a magnetic unit cell ($S_{1/2}$); (c) Thin flakes of an A-type antiferromagnet: flakes of odd layer number NN have a net magnetic moment, while those with even NN do not. (d) Two layers of the MnBi2Te4 crystal and magnetic structure, demonstrating the A-type AFM phase. Opposite spins are related by $S_{1/2}$.
@@
We investigate MnBi2Te4, a layered topological AFM with a Néel temperature of 25 K [35], as an example of a $\mathcal{T}S_{1/2}$ - symmetric AFM. Spins in individual Mn layers in MnBi2Te4 are parallel to each other, and antiparallel to the spins in neighboring layers (Fig. 1d). This order, called A-type, persists to the surface of MnBi2Te4 [40]. MOKE was recently reported in the wavelength range of $500-1000\,\mathrm{nm}$ in thin flakes of MnBi2Te4, in samples with an odd number of layers ($N$), which have a net magnetization, and in samples with an even $N$, which do not [36, 48, 39, 5]. While symmetry permits MOKE for all $N$, it does not identify the underlying mechanism.
@@
A distinct mechanism for MOKE in $\mathcal{T}S_{1/2}$ symmetric AFMs was proposed by Dzyaloshinskii and Papamichail [12]. They treated the layered AFM as a stack of FMs with alternating spin direction. The dielectric tensor of each layer is assumed to be the same as in an unbounded medium with the same FM order, and MOKE arises due to the variation in dielectric properties as the wave propagates. In the ‘alternating FM’ framework, the optical response arises from the bulk magnetic structure, and the surface conductance does not play a special role, unlike in the axion electrodynamics mechanism, where the surface contribution is essential. To the best of our knowledge, MOKE arising from the alternating FM effect has not been experimentally recognized prior to this work.
Below we report measurements of MOKE in ‘infinite’ ($\sim 136\,\mathrm{nm}$ thick, Fig. ED1) layer MnBi2Te4 as a function of magnetic field, $H$, and wavelength of the optical probe, $\lambda$. We find that non-zero MOKE appears in the $M=0$ state, and that its spectrum is quantitatively captured by the alternating FM model [12], without invoking axion electrodynamics. The alternating FM mechanism is allowed in all collinear AFMs whose spins are oriented perpendicular to the layers, regardless of the number of layers, band topology or symmetry, opening the door for MOKE studies of a broader class of AFMs than was previously considered feasible.
@@
Figure 2: Magneto-optical measurements of MnBi2Te4. (a) Reflection circular dichroism (RCD) measured at $525\,\mathrm{nm}$ as a function of magnetic field, showing a jump at the spin-flop transition ($H_{SF}\approx 3.4\,\mathrm{T}$) and a hysteresis loop with a coercive field of $H_{C}\approx 1\,\mathrm{T}$. A small setup-induced background contribution has been subtracted (Fig. ED5). RCD is compared with bulk magnetization measurements taken at $15\,\mathrm{K}$ with the field applied along the crystallographic $c$-axis (black line). The inset illustrates the bulk spin-flop transition. (b) Spatially resolved RCD measurements at $H=0$, revealing AFM domain structures that change between cooldowns.
The complex MOKE angle is defined by the relation, $\Theta\equiv-i\delta r/r$, where the reflectivities of incident LCP and RCP light are given by $r=r_{0}\pm\delta r$ (the sign and polarization conventions are defined in the Methods section). Throughout this paper we show measurements of the imaginary part of $\Theta$, which is known as reflection circular dichroism (RCD). We measure RCD using a photoelastic modulator to vary the helicity of the incident light at $50\,\mathrm{kHz}$, using a setup illustrated in Fig. ED2. The change in reflectivity synchronous with the helicity modulation is proportional to RCD (see Methods for details of the experiment).
@@
The existence of two discontinuities in RCD vs. $H$ provides an unusual opportunity to compare the MOKE spectrum in a phase with $M\neq 0$, that is the spin-flop (SF) phase, and a $\mathcal{T}S_{1/2}$ -symmetric AFM ($M=0$) phase in the same material. We measured the magnetic field dependence of RCD in the photon energy range from $1.34\,\mathrm{eV}$ ($925\,\mathrm{nm}$) to $3.22\,\mathrm{eV}$ ($385\,\mathrm{nm}$) (Fig. ED6), and extracted RCD spectra in the two phases. In Fig. 3a we compare the RCD spectrum corresponding to the SF transition with the RCD spectrum in the AFM phase. The AFM spectrum is half of the difference between RCD at $H=0^{+}$ and $H=0^{-}$, where $H=0^{+}$ and $H=0^{-}$ denote zero field as approached from positive and negative fields. The SF spectrum, corresponding to the jump in magnetization across the transition, is given by $\Theta_{SF}=\Theta(3.4\,\mathrm{T})-\left[\Theta(3.2\,\mathrm{T})-\Theta(0^{+})\right]$, because $\Theta(3.2\,\mathrm{T})$ is a sum of the antiferromagnetic and magnetization contributions, while in the spin flop phase, at $3.4\,\mathrm{T}$, there is no antiferromagnetic contribution.
@@
Figure 3: Spectroscopic signature of MOKE in antiferromagnetic MnBi2Te4. (a) Measured RCD spectrum at the spin-flop transition (purple) and at zero field (red), showing distinct spectral shapes. For each photon energy, we measured the magnetic field dependence of RCD (Fig. ED6). The AFM spectrum is half of the difference between RCD on the downward ($H=0^{+}$) and upward ($H=0^{-}$) field sweep, averaged across the hysteresis loop, between $\pm 0.9\,\mathrm{T}$. The SF spectrum is given by $\Theta_{SF}=\Theta(3.4\,\mathrm{T})-\left[\Theta(3.2\,\mathrm{T})-\Theta(0^{+})\right]$. The error bars for the $H=0^{+}$ measurement represent the standard deviation across this field range. The larger error bars for the SF spectrum reflect the fact fewer measurements were used to extract the corresponding RCD value (at $H=3.4\,\mathrm{T}$ and $H=3.2\,\mathrm{T}$). Laser power fluctuations do not contribute to these error bars, since the RCD is normalized by the simultaneously measured reflectivity (the same detector voltage is demodulated at two frequencies). The only relevant systematic uncertainty is from the PEM calibration, which for our Hinds PEM-100 is $\leq 5\%$, corresponding to <0.1<0.1 mrad on the RCD scale.
(b) RCD spectra computed from the Lorentz model (Eq. 1, with parameters from Table ED1) for a ferromagnet (purple, schematic in panel c) and two antiferromagnetic domains (red and blue, schematics in panel d). The ferromagnetic RCD is scaled by a factor of $0.4$ from that calculated for a saturated moment, to be directly comparable to the measured RCD jump at the spin-flop transition.
@@
where $f$, $\omega_{0}$, and $\gamma$ are oscillator strength, resonant frequency, and damping. These parameters, together with a background contribution, $\varepsilon_{\infty}$, are chosen to reproduce the measured optical conductivity in the spectral range of interest, reported in Ref. [27]. The good agreement between the measured quantities and those calculated from the Lorentz model is shown in Fig. ED3. Finally, $\delta\omega$, the difference in resonance frequency for LC and RC polarized light, is chosen to fit the magnitude of the measured RCD. The purple curve in Fig. 3b shows that the single oscillator model captures the main features of the observed ferromagnetic RCD spectrum (Fig. 3a).
Our quantitative description of RCD in the ferromagnetic phase of MnBi2Te4 leads us to consider the alternating FM model for the RCD in its AFM state: a stack of FM layers in which the sign of $\delta\omega$ alternates in neighboring layers. We use the transfer matrix formalism [49], following the approach in Ref. [21], to calculate the reflection coefficient from the stack. The number of layers is 100, and the thickness of each layer, $d=\mathrm{1.3}{nm}$, corresponds to the separation between spins in MnBi2Te4. The simulated RCD spectra for the two $\mathcal{T}$-reversed AFM domains (Fig. 3d), are shown in red and blue in Fig. 3b. Remarkably, this minimal model captures the spectral shape and the magnitude of RCD in the AFM phase of MnBi2Te4, without introducing free parameters beyond those that describe the FM phase. In the SI we extend this analysis by using a dielectric function computed from a density functional theory (DFT) calculation to construct the alternating FM model, and we find that this approach also yields the correct magnitude of RCD in the AFM phase (Fig. ED4). To the best of our knowledge, this is the first experimental evidence for an RCD arising from an alternating FM model in a bulk AFM.
@@
Further, the dichroism spectra for even and odd layer stacks are strikingly different (Fig. 4c, d). For $N=8$, the RCD spectrum is identical to the spectrum of the bulk AFM crystal (Fig. 3b), while spectrum for $N=9$ resembles that of the bulk SF phase, as $M\neq 0$ for odd layer stacks. The maximum RCD is about three times larger for $N=9$ than for $N=8$. However, these calculations assume that samples are suspended in vacuum, which is not a realistic experimental scenario. The transfer matrix approach allows us to include the substrate in the calculation of reflectivity [21], and we find that doing so preferentially suppresses dichroism in odd-layer samples. In Fig. ED7 we show that the maximum magnitude of RCD for samples on sapphire and diamond substrates, used in Ref. [39], is approximately equal in magnitude for even and odd layer stacks, while they maintain the characteristic spectral shapes shown in Fig. 4c. We therefore find that the alternating FM model with realistic parameters captures the main experimental results on $N$-even and $N$-odd flakes of MnBi2Te4 reported in [39], as well as on ‘infinite’ layer flakes reported here.
@@
Eq. 2 showcases the common microscopic origin of the MOKE response of an AFM bilayer and that of the corresponding ferromagnet, but it also reveals a surprising aspect of the link between them. It is natural to assume that a $\mathcal{T}S_{1/2}$ - symmetric type-A AFM can exhibit nonzero MOKE only if strong absorption (governed by the imaginary part of $n$) leads to disproportionate sampling of the top layer. However, Eq. 2 shows that this is not correct: the magnitude of $\Theta_{AFM}$ depends on the absolute value of the index, $|n|$, and is nonzero even in the absence of dissipation. The phase of $n$ appears in the phase shift between $\Theta_{AFM}$ and $\Theta_{FM}$, which manifests as the difference in the spectrum of RCD between AFM and FM structures, as seen for example in Fig. 3a.
Figure 5: Illustration of the model for MOKE in an antiferromagnetic bilayer. (a) Schematic representation of a bilayer, consisting of two layers of indices of refraction $n_{1,2}=n\pm\delta n$ when probed with circularly polarized light. $\delta n$ is induced by the opposite magnetization in the two layers, $\pm M_{z}$. (b,c) Simplified representation of a bilayer, where the layers are replaced by two infinitesimally thin sheets of conductances $G_{1,2}=G_{0}\pm\delta G$, separated by a dielectric medium of thickness $d$ and index of refraction $n$. The propagation-induced differences between fields emitted from the two sheets cancel in (b) transmission, but not in (c) reflection, leading to vanishing TCD and non-zero RCD.
@@
$E_{1,2}(z)=\frac{\mu_{0}c}{2}K_{1,2}e^{ik_{0}\left|z\pm d/2\right|},$ (4)
@@
$E_{T}(z)=\mu_{0}cE_{0}\left(\frac{G_{1}+G_{2}}{2}\right)e^{ik_{0}z}=\mu_{0}cE_{0}G_{0}e^{ik_{0}z}.$ (5)
@@
In contrast, the propagation factors for fields radiated in the z-z direction add, rather than cancel: the incident EM wave travels further to excite current in layer 2 and its radiated field propagates further before reaching an observer at z<d/2z<-d/2 (Fig. 5c). The total reflected electric field is given by,
@@ Conclusion
In this paper we identify alternating FM layers as the origin of MOKE in an A-type antiferromagnet with $\mathcal{T}S_{1/2}$ symmetry, therefore providing first experimental confirmation of the mechanism proposed by Dzyaloshinskii and Papamichail [12]. Within this theory, $\Theta_{AFM}$ is entirely determined by the MOKE response of the ferromagnetic layers that are the building blocks of the AFM, allowing us to test it directly by tuning MnBi2Te4 across the spin-flop transition, thus accessing $\Theta_{AFM}$ and $\Theta_{FM}$ in the same crystal. Based on the quantitative agreement between the measured RCD spectrum in the AFM phase and the theoretical prediction based solely on the FM spectrum, we conclude that MOKE observed at visible and near infrared wavelengths in MnBi2Te4 arises from alternating FM layers, rather than axion electrodynamics associated with band topology. Axion electrodynamics may dominate the MOKE response at terahertz frequencies below the bulk band gap [30, 1, 20, 45], where there are no other optical transitions that could overwhelm their contribution, although results in even-layer flakes suggest that the axion contribution in this frequency range is weak [30, 20].
The effect we observe is not limited to MnBi2Te4, and is generically present in A-type antiferromagnets with spins out of plane. We believe that it has not been recognized experimentally in part because the signal vanishes when averaged over atomic steps, naturally present in as-grown bulk crystals. However, optical experiments in van der Waals materials regularly probe surfaces that are atomically flat across the length scale of a laser beam. We note that some previous sightings of MOKE in van der Waals AFMs with an even number of layers were conditionally attributed to extrinsic magnetization [24, 50, 48], given the widespread conviction that MOKE intrinsic to A-type AFMs was forbidden by symmetry. While this explanation is certainly plausible, our findings show that intrinsic MOKE in A-type AFMs is always allowed, and other mechanisms should be considered when $\Theta_{AFM}$ cannot account for the magnitude, or the spectrum, of the observed signal.
@@
Our findings show that MOKE is a powerful and general tool for investigating $\mathcal{T}S_{1/2}$ symmetric AFMs. In this work we used it to image time-reversed domains, as shown for in Fig. 2b.
Further, the nonzero $\Theta_{AFM}$ enabled the discovery that a magnetic field can select antiferromagnetic domains (Fig. 2a) in such magnets, challenging the conventional view that antiferromagnets cannot be controlled by external magnetic fields. Uncovering the mechanism behind field control is beyond the scope of this study, but we note that the weakened exchange interaction at the surface is a plausible cause [40, 9]. This implies that details of the surface can affect the coercive field, as confirmed by measurements on two flakes exfoliated from the same batch (Fig. S1). However, the magnitude of the RCD at $H=0$ is a bulk property, independent of local surface conditions. The coercivity and field-switching process will be the subject of future computational and experimental investigations, enabled by the discovery of the AFM RCD signal. In summary, our results address fundamental questions concerning the origin of MOKE in AFMs, provide a strategy for studying and controlling their domains, and elucidate results on exfoliated van der Waals AFMs.